OpenStudy (anonymous):

solve the differentials:

OpenStudy (anonymous):

\[m/\sqrt{1-v^2/c^2} *(dv/dt-g)=Eq\]

OpenStudy (anonymous):

we know m=1.67*10^-27 g=9.8 E=6.2*10^5 e=1.6*10^-19 and I just need help with separating it please